OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

y2 = 2x y = 1/2 x^2 y2 = 8x y = 1/8 x^2

OpenStudy (phi):

I would plot the focus point (0,2) and the line y=-2 the parabola will open up (a smiley) that means it will be x^2 = 4 p y where p is 1/2 the distance between the focus and the directrix here p is 2 so the equation will be x^2 = 8 y